OpenStudy (anonymous):

Cool treats sold 60 ice cream cones.Single dip cones sold for $2.50 each and double dip cones for $4.15 each.In all $179.70 was taken in for the cones.How many of each size cone were sold?.I was helped for most of it but am stuck at the multiplying part.

OpenStudy (calculusxy):

Use the system of equations.

OpenStudy (calculusxy):

x = single dip y = double dip 60 = total ice creams x + y = 60 our second equation would be: $2.50x + $4.15y = $179.70 you can use the substitution or the elimination method to solve for either x or y.

OpenStudy (calculusxy):

Are you familiar with this concept?

OpenStudy (anonymous):

I am just now grasping the method . I am stuck at the part of solving for S and D

OpenStudy (calculusxy):

Let's solve for x = single dip We will take the easy equation: x + y = 60 subtract x from both sides y = 60 - x now your value for y (in the second equation) is y = 60 -x Now place it into the second equation: $2.50x + $4.15(60 - x) = $179.70 this is so much easier to do because now you have only one variable to solve for (x), instead of two ( x and y ). solve for x and we will move on to the next step.

OpenStudy (calculusxy):

Do you understand what I just did?

OpenStudy (calculusxy):

If you don't, then you could ask me what part you're confused on.

OpenStudy (anonymous):

The part where I am subtracting the 60 from x?

OpenStudy (calculusxy):

We want to solve for x, and that will lead us to finding the amount of single dip cones. The reason for why we are isolating y is to get the value of it. Instead of having two variables in front of us, we will have one (and that is x -- the variable that we are solving for).

OpenStudy (calculusxy):

Therefore, we can plug in y = 60 - x into the second equation where y is (next to $4.15). if you look at that then you can see that there is only x.